### v1

# Quantum reservoir computing maps data onto the Krylov space

Abstract
We demonstrate that time-evolved operators can construct a Krylov space to compute Operator complexity and introduce Krylov observability as a measure of phase space dimension in quantum systems. We test Krylov observability in the framework of quantum reservoir computing and show that it closely mirrors information processing capacity, a data-driven expressivity metric, while achieving computation times that are orders of magnitude faster. Our results validate Operator complexity and give the interpretation that data in a quantum reservoir is mapped onto the Krylov space.
preprint: APS/123-QED
Introduction. — The significance of physics in machine learning, and vice versa, is unmistakable, as evidenced by the awarding of the Nobel Prize to John Hopfield and Geoffrey Hinton. An example of the intersection of physics and machine learning is quantum computing, a technology that seeks to harness quantum effects for computation. Key research questions in quantum computation are the extent to which machine learning algorithms can be implemented on quantum systems, how they might function, and what advantages this could yield. Therefore, understanding quantum machine learning networks is of utmost importance. In classical machine learning, understanding is achieved through methods of expressivity and explainability. In quantum mechanics, quantum information measures are used to gain deeper insights. This raises the question of how advancements in quantum information can be leveraged to enhance our understanding of quantum machine learning.
A promising information measure in this context is Krylov complexity, which quantifies how much the evolved operators or states spread within the Krylov basis [1, 2]. It has been shown that Operator complexity measures provide insights into quantum chaos[3], spin chains, SYK models[2, 4, 5], 2D CFT [6], closed systems, driven systems [7], and open quantum systems [8], among many others [9, 10, 11, 2, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40].
Operator complexity can be interpreted as the expectation value of the associated Krylov operator. We believe that a measure on the Krylov phase space can offer further insights and propose Krylov observability. Unlike Operator complexity, Krylov observability accommodates multiple operators, respects an upper bound defined by the phase space dimension, and integrates multiple measurements, which are all necessities for quantum machine learning. The defined spaces can also be extended to define a Operator complexity measure, which effectively consider multiple operators.
We test this measure in the context of quantum reservoir computing. Unlike variational quantum machine learning, quantum reservoir computing utilizes a physical system in which only the readout layer is trained [41, 42, 43, 44, 45]. Extensive research has been conducted to optimize quantum reservoir computing, ranging from leveraging noise and decoherence, improving task performance and developing various methods to address the time-complexity problem [45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64]. An established measure of expressivity in reservoir computing is the information processing capacity ($\mathrm{IPC}$) [65], which quantifies the reservoir’s ability to map and retain data both linearly and non-linearly.
In this work we demonstrate that time-evolved operators can construct a Krylov space to compute Operator complexity and introduce Krylov observability as a measure of the phase space dimension. We then compute Krylov observability for various systems and observables, and compare the behavior with the information processing capacity. We demonstrate that Krylov observability and information processing capacity exhibit correlation factors as high as 0.97. Our results show that quantum reservoirs map data onto the Krylov space and that sampling this space can be indicative of task performance. Our findings demonstrate the utility of Krylov observability as a measure of the computational capacity of quantum systems which is both upper bounded and computationally inexpensive (calculating the information processing capacity for the results in Figure 1 took 150 hours per reservoir, whereas Krylov observability was computed in just 30 seconds per reservoir).
Spread complexity was first discussed in [66] in relation to quantum reservoir computing, where the authors show that spread complexity can be used to explain task performance. In [40], we perform a thorough study of various Krylov-based measures in relation to understanding task performance in QRC. We discuss Krylov observability, which is introduced in this work, as well as fidelity, spread complexity, Krylov expressivity (introduced in [40]), and the information processing capacity in the context as expressivity measures for quantum reservoir computers. There, we show that more expressive quantum systems can exhibit worse task performance if the considered observables commute with a larger subspace of the system.
Krylov spaces — The evolution of an operator $O$ with a Hamiltonian $H$ is given by $\partial_{t}O(t)=i[H,O(t)]$. The solution to this equation is
$O(t)$ $=e^{iHt}Oe^{-iHt}=\sum_{k=0}^{\infty}\frac{(it)^{k}}{k!}\mathcal{L}^{k}(O),$
where $\mathcal{L}(O)=HO-OH$ represents the Liouvillian. This implies that $O(t)\in\mathrm{Span}\{\mathcal{L}^{0}(O),\mathcal{L}^{1}(O),\ldots\}$. The authors in [1] used the Krylov space property and the linearity of $\mathcal{L}(O)$ to show that there exists an $M\in\mathbb{N}$ such that
$O(t)\in\mathrm{Span}\{\mathcal{L}^{0}(O),\mathcal{L}^{1}(O),\dots,\mathcal{L}^{M-1}(O)\}=\mathcal{L}_{M}$
holds true. Orthonormalizing $\mathcal{L}_{M}$ results in the basis $\{\mathcal{W}_{i}\}_{i}$, such that the space can be rewritten as $\mathcal{L}_{M}=\mathrm{Span}\{\mathcal{W}_{0},\mathcal{W}_{1},\dots,\mathcal{W}_{M-1}\}$. With the representation of the time-evolved operator $O(t)$ in the Krylov basis, the Operator complexity $\mathcal{C}_{O}(t)$ is defined as follows [1]:
$O(t)$ $=\sum_{n=0}^{M-1}i^{n}\beta_{n}(t)\mathcal{W}_{n},$
$\mathcal{C}_{O}(t)$ $=\sum_{n=0}^{M-1}(n+1)\absolutevalue{\beta_{n}(t)}^{2}.$ (1)
Krylov Observability— In this work, we will prove that time-evolved observables can be used to construct a space that is equivalent to the space $\mathcal{L}_{M}$. Building upon these results, we will define an observability measure to predict task performance and further our understanding of quantum reservoir computing.
Theorem 1.
Let $O\in\mathbb{C}^{N,N}$ be a Hermitian observable, i.e., $O=O^{\dagger}$. Let $O(t)=e^{iHt}Oe^{-iHt}$, where $H$ is the Hamiltonian, and let $M$ be the Krylov grade, such that $O(t)\in\mathcal{L}_{M}$. Let $0=t_{0}<t_{1}<t_{2}<\dots<t_{Q}$ and define the approximate evolved operator as
$\tilde{O}(t_{a})=\sum_{k=0}^{Q}\frac{(it_{a})^{k}}{k!}\mathcal{L}^{k}(O).$
Then, it holds that $\mathcal{L}_{Q}=\tilde{\mathcal{F}}_{Q}$, for all $Q\in\mathbb{N}$, where ~$\tilde{\mathcal{F}}_{Q}$ is given by
~$\tilde{\mathcal{F}}_{Q}=\mathrm{Span}\{\tilde{O}(t_{0}),\tilde{O}(t_{1}),\dots,\tilde{O}(t_{Q})\}.$
In the limit as $Q\rightarrow\infty$, it follows that $\mathcal{L}_{M}=\mathcal{F}_{M}$, where $\mathcal{F}_{Q}$ is given by
${\mathcal{F}}_{Q}=\mathrm{Span}\{{O}(t_{0}),{O}(t_{1}),\dots,{O}(t_{M-1})\}.$
Proof.
$\tilde{O}(t_{a})$ can be represented as
$\tilde{O}(t_{a})$ $=(i^{0}\mathcal{L}^{0}(O),i^{1}\mathcal{L}^{1}(O),\dots,i^{Q}\mathcal{L}^{Q}(O))\begin{pmatrix}1\\ t_{a}\\ \vdots\\ \frac{t_{a}^{Q}}{Q!}\end{pmatrix}$
Representing $Q+1$ time-evolved $\tilde{O}(t_{a})$ with pairwise distinct times $t_{0}<t_{1}<\dots<t_{Q}$ leads to
$(\tilde{O}(t_{0}),\dots,\tilde{O}(t_{Q}))=(i^{0}\mathcal{L}^{0}(O),\dots,i^{Q}\mathcal{L}^{Q}(O))\Theta,$
$\mathrm{with~{}}$ $\Theta=\begin{pmatrix}1&1&1&\dots&1\\ {t_{0}}/{1!}&{t_{1}}/{1!}&{t_{2}}/{1!}&\dots&{t_{Q}}/{1!}\\ {t_{0}^{2}}/{2!}&{t_{1}^{2}}/{2!}&{t_{2}^{2}}/{2!}&\dots&{t_{Q}^{2}}/{2!}\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ {t_{0}^{Q}}/{Q!}&{t_{1}^{Q}}/{Q!}&{t_{2}^{Q}}/{Q!}&\dots&{t_{Q}^{Q}}/{Q!}\end{pmatrix}.$
The pairwise distinct times $t_{a}$ imply the linear independence of each column and the invertibility of $\Theta$, resulting in
$(\tilde{O}(t_{0}),\dots,\tilde{O}(t_{Q}))\Theta^{-1}=(i^{0}\mathcal{L}^{0}(O),\dots,i^{Q}\mathcal{L}^{Q}(O)).$
Since all vectors of $\mathcal{L}_{Q}$ can be represented through vectors in ~$\tilde{\mathcal{F}}_{Q}$, it follows that $\mathcal{L}_{Q}=\tilde{\mathcal{F}}_{Q}$ holds for all $Q$. In the limit as $Q$ approaches infinity, $\tilde{O}(t_{a})\rightarrow O(t_{a})$ holds, such that the space ~$\tilde{\mathcal{F}}_{Q}$ can be expressed as:
$\mathcal{L}_{Q}=\tilde{\mathcal{F}}_{Q}=\mathrm{Span}(O(t_{0}),O(t_{1}),\dots,O(t_{Q}))={\mathcal{F}}_{Q}.$ (2)
For $Q>M$, $\mathcal{L}_{Q}$ has only $M$ linearly independent vectors, implying that $\mathcal{F}_{Q}$ has only $M$ linearly independent vectors, thus resulting in $\mathcal{L}_{M}=\mathcal{F}_{M}$. ∎
In quantum reservoir computing, the expectation values of $K$ observables $O_{1},O_{2},\dots,O_{K}$ are sampled at discrete times $\expectationvalue{O_{i}(t_{j})}$. From Theorem 1
$O_{i}(t)\in\mathcal{F}_{i}=\mathrm{Span}(O_{i}(t_{0}),O_{i}(t_{1}),\dots,O_{i}(t_{M_{i-1}}))$ (3)
follows. $M_{i}$ is the dimension of the space, and $t_{M_{i-1}}=T_{O}$ is equal for all observables $O_{i}$. For any $O_{i}\in\{O_{1},O_{2},\ldots,O_{K}\}$, it holds that
$O_{i}(t)\in\bigcup_{i=1}^{K}\mathcal{F}_{i}=:\mathcal{F}(O_{1},O_{2},\dots,O_{K}),\quad\forall O_{i}.$ (4)
In the following, $\mathcal{F}$ always denotes the full space, i.e., $\mathcal{F}:=\mathcal{F}(O_{1},O_{2},\dots,O_{K})$. Constructing the spaces in this way does not consider the linear independence of the spaces $\mathcal{F}_{k}$ and does not provide information about how much each $\mathcal{F}_{k}$ contributes to $\mathcal{F}$. Algorithm 1 constructs spaces where both factors are considered. Let $\mathcal{F}^{B}$ and $\mathcal{F}^{B}_{i}$ be the bases of the spaces $\mathcal{F}$ and $\mathcal{F}_{i}$, i.e., $\mathcal{F}=\mathrm{Span}(\mathcal{F}^{B})$ and $\mathcal{F}_{i}=\mathrm{Span}(\mathcal{F}^{B}_{i})$.
Algorithm 1 Construction of Observability Spaces
1:$\mathcal{I}_{O}=\{1,2,\dots,K\}$
2:$\mathcal{T}=\{t_{1},t_{2},\dots,t_{R}\}$
3:$\mathcal{F}^{(B)}=\emptyset$
4:~$\tilde{\mathcal{F}}_{1}^{(B)},\tilde{\mathcal{F}}_{2}^{(B)},\dots,\tilde{\mathcal{F}}_{K}^{(B)}=\emptyset,\emptyset,\dots,\emptyset$
5:while $t_{j}\in\mathcal{T}$ do
6:     for $k\in\mathcal{I}_{O}$ do
7:         if $O_{k}(t_{j})\notin\mathcal{F}$ then
8:              $\mathcal{F}^{(B)}=\mathcal{F}^{(B)}\cup O_{k}(t_{j})$
9:              ~$\tilde{\mathcal{F}}_{k}^{(B)}=\tilde{\mathcal{F}}_{k}^{(B)}\cup O_{k}(t_{j})$
10:         end if
11:     end for
12:end while
13:$\mathcal{F}^{(B)}\leftarrow\mathrm{orthonormalize}(\mathcal{F}^{(B)})$
In the first two lines, the index set $\mathcal{I}_{O}$ for the observables and the set of all times $\mathcal{T}$ are initialized, where $R>M$ should hold, and in our case, $R=\mathrm{dim}(H)$. Lines 3 and 4 initialize all bases to the empty set, and line 5 iterates through all times $t_{j}$. The second while loop in line 6 iterates through all observables $O_{k}$. If the $k$-th time-evolved observable $O_{k}(t_{j})$ is not in the space $\mathcal{F}=\mathrm{Span}(\mathcal{F}^{B})$, then the time-evolved observable is added to the bases $\mathcal{F}^{(B)}$ and ~$\tilde{\mathcal{F}}_{k}^{(B)}$ in lines 8 and 9. This can be verified by computing the rank of the basis and checking whether it changes. If the rank does not change, i.e., $\mathrm{rank}(\mathcal{F}^{(B)})=\mathrm{rank}(\mathcal{F}^{(B)}\cup O_{k}(t_{j}))$, then it follows that $O_{k}(t_{i})\in\mathcal{F}$ holds as well. In line 13, the basis $\mathcal{F}^{(B)}$ is orthonormalized, such that spaces with the following properties are returned:
$\mathcal{F}=\bigcup_{k=1}^{K}\mathcal{F}_{k}=\bigcup_{k=1}^{K}\tilde{\mathcal{F}}_{k}~{},~{}~{}\mathrm{dim}\left(\bigcup_{j=1}^{l}\tilde{\mathcal{F}}_{j}\right)=\sum_{j=1}^{l}\mathrm{dim}(\tilde{\mathcal{F}}_{j})$
This results in the spaces $\mathcal{F}_{k}$ being linearly independent and ensures that each observable $O_{k}$ is mapped onto a space $\mathcal{F}_{k}$. In the next step, we use these spaces to define Krylov observability, a measure to capture the phase space dimension of various operators.
Proposition 1 (Krylov Observability).
Assume $K$ observables $O_{1},\dots,O_{K}$. For each observable $O_{k}$, the spaces ${\mathcal{F}}_{k}$ are computed with $\mathrm{dim}({\mathcal{F}}_{k})=M_{k}$ and let $T>0$. With $R_{k}=\mathrm{min}(V,M_{k})$ and $\tau_{k}=T/R_{k}$, define
$\mathcal{G}_{k}=\{O_{k}(\tau_{1}),O_{k}(\tau_{2}),\dots,O_{k}(\tau_{M_{k}})\}$
where $V$ is the number of measurements for each observable. Upper bounding the number of times $R_{k}$ by $M_{k}$ ensures that $\mathcal{G}_{k}$ always consists of the smallest number of linearly independent basis states. The observability of the $k$-th observable $O_{k}$ is defined as
$p_{k}(T)=1+\sum_{j=1}^{R_{k}-1}(1-f(O_{k}(\tau_{j}),O_{k}(\tau_{j+1}))),$ (5)
where $f$ is a function with $f(A,B)=1$ if $A=B$. In this study, we set
$|\displaystyle f(A,B):=\mathrm{F}(A,B)=\absolutevalue{\mathrm{Tr}(\frac{A^{\dagger}}{\norm{A}}\frac{B}{\norm{B}}}$ | (6)
$\mathrm{F}$ is the normalized fidelity between the two time-evolved operators. The Krylov  Observability $\mathcal{O}_{K}(T)$ of $V$ multiplexed observables $O_{1},\dots,O_{K}$ is defined as
$\mathcal{O}_{K}(T)=\sum_{k=1}^{K}p_{k}(T).$ (7)
Quantum Reservoir Computing— We demonstrate the utility of Krylov Observability using the widely researched Ising Hamiltonian
$H_{I}=\sum_{i=1,j>i}^{N_{S}}J_{ij}X_{i}X_{j}+\sum_{i=1}^{N_{S}}hZ_{i}.$ (8)
as a quantum reservoir. $X_{i}$ and $Z_{i}$ represent the Pauli-$x$ and Pauli-$z$ operators acting on the $i$-th site, and $N_{S}$ is the number of sites. The constant $h$ is set to $h=0.5$, and $J_{i,j}$ is sampled from a uniform distribution on the interval $[0.25,0.75]$[45]. The system evolution under the von Neumann equation is given by
$\partial_{t}\rho=-i[H,\rho].$ (9)
[…]
Reservoir computing research often makes use of the information processing capacity ($\mathrm{IPC}$) as an expressivity measure. The $\mathrm{IPC}$ evaluates how well prior inputs and combinations of prior inputs can be reconstructed non-linearly [65]. The first-order $\mathrm{IPC}_{1}$, also called memory capacity, measures how well the system can reconstruct prior data [67]. The definition for computing the $\mathrm{IPC}$ is given in [65, 61, 68, 69, 70]. The sum over all $\mathrm{IPC}$s is called the total $\mathrm{IPC}$, or simply $\mathrm{IPC}$, i.e., $\mathrm{IPC}=\sum_{i}\mathrm{IPC}_{i}$. [71, 72] provides analytical evidence of the influence of eigenvalues on the $\mathrm{IPC}$. Furthermore, it has been shown that $\mathrm{IPC}$ can help predict task performance in reservoir computing [73, 74], and there is a typically observed non-linearity-memory trade-off, which was observed in quantum reservoirs as well [61].
Results—To analyze the ability of our Krylov observability measure $\mathcal{O}_{K}$ to quantify the computational capabilities of a quantum system, we compare it with the $\mathrm{IPC}$. We consider ten different Ising Hamiltonians, each with four sites and uniformly distributed coupling constants $J_{i,j}$ on the interval $[0.25,0.75]$. For each Hamiltonian, we consider a quantum reservoir where the first site ($Z_{1}$), the first two sites ($Z_{1,2}$), and all four sites ($Z_{1,2,3,4}$) are measured in the Pauli-$z$ direction. We compute the information processing capacity $\mathrm{IPC}$ and the Krylov observability $\mathcal{O}_{K}$ for various clock cycle lengths $T$ and for various time-multiplexing measurements $V$.
Figure 1: Observability $\mathcal{O}_{K}$ and $\mathrm{IPC}$ in dependence on multiplexing $V$ and clock cycle $T$.
Figure 1 shows the results for the averaged $\mathrm{IPC}$ (left column) and the averaged $\mathcal{O}_{K}$ (right column). In each plot, the $y$-axis represents the number of measurements for the corresponding measurement scheme and the $x$-axis represents the clock cycle $T$. The first row shows results when only the first site is measured $Z_{1}$.
As the clock cycle $T$ and the number of measurements $V$ increase, so does the phase space dimension, i.e., the Krylov observability $\mathcal{O}_{K}$, onto which the data is mapped. At $T_{\mathrm{sat}}(Z_{1})\approx 36$ and $V_{\mathrm{sat}}(Z_{1})\approx 110$, the full Krylov space is explored, as indicated by the saturation of $\mathcal{O}_{K}$. Further increases in the number of measurements and clock cycles result only in measurements that are linearly dependent on all previous ones, thus not further increasing the expressivity of the reservoir. This is observed in the almost identical behavior of $\mathrm{IPC}$ and $\mathcal{O}_{K}$. To better quantify the behavior of the Krylov observability $\mathcal{O}_{K}$ and information processing capacity $\mathrm{IPC}$, we compute the Pearson correlation factor between them. In the case where only one observable is measured, a Pearson correlation factor of $P_{C}=0.97$ is reached. Measuring the first two sites yields results in the second row, where both measures show similar behavior with slightly smaller values for $\mathcal{O}_{K}$. The number of virtual nodes at which $\mathrm{IPC}$ and $\mathcal{O}_{K}$ saturate reduces to $T_{\mathrm{sat}}(Z_{1,2})\approx 30$ and $V_{\mathrm{sat}}(Z_{1,2})\approx 95$ with $P_{C}=0.97$. Measuring all four sites moves the saturation to smaller values, $V_{\mathrm{sat}}(Z_{1,2,3,4})\approx 75$ and $T_{\mathrm{sat}}(Z_{1,2})\approx 18$. The earlier saturation when measuring multiple observables is due to the higher number of measurements $N_{R}$ as the total number of measurements is given by $N_{R}=VK$. Additionally, the space dimension of $\mathcal{F}(Z_{1})$ is smaller than that of $\mathcal{F}(Z_{1},Z_{2})$ and $\mathcal{F}(Z_{1},Z_{2},Z_{3},Z_{4})$. One interesting observation, particularly for $\mathcal{O}_{K}$ in the third row, is oscillations over time, arising from the observability measure construction, where the fidelity of time-evolved observables is computed. To better understand this behavior, Fig. 2 shows the $\mathrm{IPC}$ (blue) and $\mathcal{O}_{K}$ (red) as functions of the clock cycle $T$ for a fixed number of measurements $V=30$ (left), and as functions of $V$ for a fixed clock cycle $T=20$ (right). In the left plot, $\mathcal{O}_{K}$ increases slightly later but reaches a maximum with oscillations, while the $\mathrm{IPC}$ reaches a maximum around $T\approx 20$ and remains stable. The standard deviation is shown as a fill in the corresponding color and is small for both systems. The oscillation is due to the small number of qubits, as quasiperiodic behavior is more apparent [40]. The right plot shows the dependence on the number of measurements $V$ for constant $T=20$. The behavior of $\mathcal{O}_{K}$ and $\mathrm{IPC}$ is, except for an offset, almost identical.
Figure 2: Observability $\mathcal{O}_{K}$ and $\mathrm{IPC}$ in dependence on multiplexing $V$ and clock cycle $T$.
Conclusion and Discussion— This work showed that instead of computing different powers of the Liouvillian, time-evolved operators can be used to construct an equivalent space in Theorem 1. Building upon this, we have shown how observability spaces can be constructed that consider the possible dependence of both spaces, and an algorithm for such a computation is provided. This enabled our work to propose Krylov observability as a measure to capture the phase space dimension of evolved observables, see Proposition 1. Krylov observability was then tested in the framework of quantum reservoir computing. We showed that the information processing capacity, consisting of thousands of quantum reservoir tasks, and the Krylov observability exhibit almost identical behavior, with correlation factors of $P_{C}=0.97$. The high correlation between the information processing capacity and Krylov observability suggests that quantum reservoir computing can be interpreted as a map of classical data onto the Krylov spaces discussed in this work. Second, the field of Krylov complexity is validated as an effective method for understanding quantum systems through Krylov observability, which can predict how well macroscopic information (i.e., encoded input series) can be reconstructed and non-linearly transformed in a quantum system. This study further demonstrates that information processing capacity, or macroscopic information, might be an interesting technique to describe the physics of quantum systems. Therefore, quantum reservoir computing might become a tool to further the understanding of quantum systems. Lastly, we show that Krylov observability serves as an expressivity measure that is fast to compute, with a simulation time for one Hamiltonian of 30 seconds, while the corresponding computation of the information processing capacity can take over 150 hours. Programming information processing capacity also requires careful consideration of combinatorics and suitable cutoffs, whereas programming Krylov observability is straightforward. Lastly, Krylov observability can easily adapt as an expressivity measure for variational quantum machine learning and quantum extreme learning.